OpenStudy (katie327):

Write the logarithmic expression as a single log with coeffcient 1, and simplify as much as possible. assume variables represent positive real numbers

OpenStudy (katie327):

\[3\log_{3} m-8\log_{3} n-2\log_{3} P\]

OpenStudy (katie327):

I have two more I need help with. I have done 47 on my own but these 3 have me stumped. \[Q=Q _{0}e ^{-kt} (used for chemistry) k=?\]

OpenStudy (katie327):

\[\ln (\frac{ k }{ A })=\frac{ -E }{ RT } k =?\]

OpenStudy (unklerhaukus):

\[c\log_ba = \log_ba^c\]

OpenStudy (unklerhaukus):

\[\log_b x+\log_b y-\log_b z=\log_b\frac{xy}{z}\]

OpenStudy (unklerhaukus):

can you simplify now?

OpenStudy (katie327):

\[\log_{3} \frac{ (m-8(n-2)}{ P }\]

OpenStudy (katie327):

?

OpenStudy (katie327):

Sorry, I am in panic mode, I only have 20 minutes left to submit and I am having bad anxiety. Thanks os much for your help

OpenStudy (unklerhaukus):

\[3\log_{3} m-8\log_{3} n-2\log_{3} P\] First we use \(\boxed{c\log_ba = \log_ba^c}\) \[\log_{3} m^3-\log_{3} n^8-\log_{3} P^2\]

OpenStudy (unklerhaukus):

Now we use \(\boxed{\log_b x-\log_b y=\log_b\frac xy}\)

OpenStudy (unklerhaukus):

What do you get?

OpenStudy (katie327):

\[\log_{3} \frac{ m ^{3} }{ n ^{8} }\]

OpenStudy (unklerhaukus):

that's right, but don't for get the P term as well

OpenStudy (katie327):

does that go on the bottom? like this? \[\log_{3} \frac{ m ^{3} }{ n ^{8}p}\]

OpenStudy (unklerhaukus):

yeah it does go there on the bottom, but it should still have the ^2 with it.

OpenStudy (katie327):

Like this? \[\log_{3} \frac{ m ^{3} }{ n ^{8}P ^{2} }\]

OpenStudy (unklerhaukus):

Yes, that's it! \(\color{red}\checkmark\)

OpenStudy (katie327):

Thank you so much !! Could you help me with the other two real quick? please?

OpenStudy (unklerhaukus):

I can try to, What is the question that goes with the chem. equation?

OpenStudy (katie327):

Thats the full question

OpenStudy (unklerhaukus):

oh, it's the same question as the first ?

OpenStudy (katie327):

its the \[\ln (\frac{ k }{ A }=\frac{ -E }{ RT }\]

OpenStudy (katie327):

and its asking what k =

OpenStudy (katie327):

I'm not sure how to do that ? :(

OpenStudy (unklerhaukus):

oh wait, did we skip one?

OpenStudy (katie327):

Yeah :) theres one above the Q one

OpenStudy (unklerhaukus):

\[Q = Q_0e^{-kt}\\ \frac{Q}{Q_0}=e^{-kt}\\ \ln\frac{Q}{Q _{0}}=\ln e ^{-kt}\\ \qquad\quad =-kt\]

OpenStudy (katie327):

I'm sorry, I just don't understand how to solve for k

OpenStudy (unklerhaukus):

\[\ln \frac{ k }{ A } = \frac{ -E }{ RT }\\ \frac kA = e^{-E/RT}\\ k=Ae^{-E/RT}\]

OpenStudy (unklerhaukus):

does this help?

OpenStudy (katie327):

Okay yes, except i still cant figure out what k= for the Q one is

OpenStudy (unklerhaukus):

\[\ln\frac Q{Q_0} =−kt\\ -\ln\frac Q{Q_0}=\ln\frac{Q_0}Q =kt\\ k = \frac1t\ln\frac{Q_0}Q\]

OpenStudy (katie327):

Oh, Okay that makes sense. I really appreciate the help!!

OpenStudy (unklerhaukus):

\[\ddot\smile \]